OpenStudy (anonymous):

What is the simplified form of 3 over 4x plus 3 + 21 over 8 x squared minus 14x minus 15 ? 6 time the quantity 2 x plus 5 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 over the quantity 4 x plus 3 6 time the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 time the quantity x plus 1 end quantity over the quantity 2x plus 5 end quantity times 4 x plus 3

OpenStudy (anonymous):

If you could retype this with either parenthesis or using the equation button below please. I would need to see that in order to determine our order of operation.

OpenStudy (anonymous):

\[\frac{ 3 }{ 4x+3}+\frac{ 21 }{ 8x^2-14x-15 }\] like this?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

can you factor the second denominator? that's the toughest part i think

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

that's a tough one to factor eh?

OpenStudy (anonymous):

i think I'll just factor it for you then help you set up the rest.

OpenStudy (anonymous):

\[\frac{ 3 }{ 4x+3 }+\frac{ 21 }{ (4x+3)(2x-5) }\] does that help?

OpenStudy (anonymous):

the first fraction needs to be multipled by (2x-5)/(2x-5) so you have a common denominator

OpenStudy (anonymous):

\[\frac{ 3(2x-5) }{ (2x-5)(4x+3) }+\frac{ 21 }{ (2x-5)(4x+3) }=?\]

OpenStudy (anonymous):

can you combine the numerators?